OpenStudy (anonymous):

simplify: 4 - 2 --- --- x-y y-x a)-6 ---- x-y b)-2 ---- x-y c) 2 ---- x-y d) 6 ---- x-y e) 6 ---- x^2-y^2

jimthompson5910 (jim_thompson5910):

Sorry, for some reason the equation editor is not working. The answer comes out to D) Here's my work (it looks better with the equation editor working) \[\frac{4}{x-y}-\frac{2}{y-x}\] \[\frac{4}{x-y}-\frac{2}{-1(-y+x)}\] \[\frac{4}{x-y}-\left(-\frac{2}{x-y}\right)\] \[\frac{4}{x-y}+\frac{2}{x-y}\] \[\frac{4+2}{x-y}\] \[\frac{6}{x-y}\] So \[\frac{4}{x-y}-\frac{2}{y-x}=\frac{6}{x-y}\] So the answer is choice D)

OpenStudy (anonymous):

thanks